# Incorporating Priors in learning: A Random matrix study under a Teacher–Student Framework

### v1

@@ 2 ASSUMPTIONS AND SETUP
${\mathbf{X}}=\left[{\mathbf{x}}_{1}^{\top};\dots;{\mathbf{x}}_{n}^{\top}\right]\in\mathbb{R}^{n\times d},\quad{\mathbf{Y}}=\left[{\mathbf{y}}_{1}^{\top};\dots;{\mathbf{y}}_{n}^{\top}\right]\in\mathbb{R}^{n\times q}.$
@@
$p({\mathbf{Y}}\mid{\mathbf{X}},{\mathbf{\Theta}})$ $\sim\prod_{i=1}^{n}\mathcal{N}({\mathbf{y}}_{i}\mid{\mathbf{\Theta}}^{\top}{\mathbf{x}}_{i},\sigma^{2}{\mathbf{I}}_{q}),$
@@
$=\left(\frac{1}{n}{\mathbf{X}}^{\top}{\mathbf{X}}+\lambda{\mathbf{I}}_{d}\right)^{-1}\left(\frac{1}{n}{\mathbf{X}}^{\top}{\mathbf{Y}}+\lambda{\mathbf{\Theta}}_{0}\right),$
@@ 3 Main Results
We sketch the proof. The uniqueness of $\delta$ follows from [17, Proposition 3.12]). Rearranging (2), using a basic trace property and taking the expectation with respect to $\mathbf{x}$},
@@
we substitute ${\mathbf{\Theta}}_{\operatorname{MAP}}={\mathbf{Q}}\left(\tfrac{1}{n}\,{\mathbf{X}}^{\top}{\mathbf{Y}}+\lambda{\mathbf{\Theta}}_{0}\right)$, where the matrix ${\mathbf{Q}}=\left(\frac{1}{n}{\mathbf{X}}{\mathbf{X}}^{\top}+\lambda{\mathbf{I}}\right)^{-1}$ is the so-called resolvent of the matrix $\frac{1}{n}{\mathbf{X}}{\mathbf{X}}^{\top}$. Expand the traces using cyclicity and using the definition of $\mathbf{Y}$} as in (1), and take expectations over the Gaussian noise contained in $\mathbf{Y}$}. This leaves only terms in $\mathbf{\Sigma}$}, $\mathbf{Q}$}, and ${\mathbf{Q}}{\mathbf{\Sigma}}{\mathbf{Q}}$. By standard random matrix arguments [17], we replace $\mathbf{Q}$} and ${\mathbf{Q}}{\mathbf{\Sigma}}{\mathbf{Q}}$ by their deterministic equivalents [16, Definition 4], given as follows by (see also [15, 18, 17])
@@ 4 Special case of Identity Covariance
Figure 1 compares theoretical and empirical risks for linear regression with Gaussian features. Solid lines show theory, dashed lines with markers show empirical means over 20 Monte Carlo runs, with error bars for one standard deviation.

### v2

@@ 2 ASSUMPTIONS AND SETUP
${\mathbf{X}}=\left[{\mathbf{x}}_{1}^{\top};\dots;{\mathbf{x}}_{n}^{\top}\right]\in\mathbb{R}^{d\times n},\quad{\mathbf{Y}}=\left[{\mathbf{y}}_{1}^{\top};\dots;{\mathbf{y}}_{n}^{\top}\right]\in\mathbb{R}^{q\times n}.$
@@
$p({\mathbf{Y}}\mid{\mathbf{X}},{\mathbf{\Theta}})$ $\sim\prod_{i=1}^{n}\mathcal{N}({\mathbf{y}}_{i}\mid{\mathbf{\Theta}}^{\top}{\mathbf{x}}_{i},\tilde{\sigma}^{2}{\mathbf{I}}_{q}),$
@@
$=\left(\frac{1}{n}{\mathbf{X}}{\mathbf{X}}^{\top}+\lambda{\mathbf{I}}_{d}\right)^{-1}\left(\frac{1}{n}{\mathbf{X}}{\mathbf{Y}}^{\top}+\lambda{\mathbf{\Theta}}_{0}\right),$
@@ 3 Main Results
We sketch the proof. The uniqueness of $\delta$ follows from [13, Proposition 3.12]). Rearranging (2), using a basic trace property and taking the expectation with respect to $\mathbf{x}$},
@@
we substitute ${\mathbf{\Theta}}_{\operatorname{MAP}}={\mathbf{Q}}\left(\tfrac{1}{n}\,{\mathbf{X}}{\mathbf{Y}}^{\top}+\lambda{\mathbf{\Theta}}_{0}\right)$, where the matrix ${\mathbf{Q}}=\left(\frac{1}{n}{\mathbf{X}}{\mathbf{X}}^{\top}+\lambda{\mathbf{I}}\right)^{-1}$ is the so-called resolvent of the matrix $\frac{1}{n}{\mathbf{X}}{\mathbf{X}}^{\top}$. Expand the traces using cyclicity and using the definition of $\mathbf{Y}$} as in (1), and take expectations over the Gaussian noise contained in $\mathbf{Y}$}. This leaves only terms in $\mathbf{\Sigma}$}, $\mathbf{Q}$}, and ${\mathbf{Q}}{\mathbf{\Sigma}}{\mathbf{Q}}$. By standard random matrix arguments [13], we replace $\mathbf{Q}$} and ${\mathbf{Q}}{\mathbf{\Sigma}}{\mathbf{Q}}$ by their deterministic equivalents [6, Definition 4], given as follows by (see also [18, 19, 13])
@@ 4 Special case of Identity Covariance
Figure1  [^1] Reproducibility code available at https://github.com/maliktiomoko/Priors_learning compares theoretical (solid) and empirical (dashed) risks. Blue/green: $\Theta_{0}\neq 0$, orange/red: $\Theta_{0}=0$, with shaded areas for std.